OpenStudy (anonymous):

Determine an equation for the line perpendicular to -3x+15y=13 passing through the points (0,3)

OpenStudy (anonymous):

what is the slope for the perpendicular line? first write your given into format y = mx + b so that you can easily see the slope of that line..

OpenStudy (anonymous):

once you determine the perpendicular slope, use point slope form to determine your new equation y - y1 = m (x - x1) given point is (x1, y1)

OpenStudy (anonymous):

y=3/15x+13/15

OpenStudy (anonymous):

great....you see the slope for the given line; what is the slope of perpendicular line?

OpenStudy (anonymous):

y-3=3/15(x-0)

OpenStudy (anonymous):

y=3/15x+3

OpenStudy (anonymous):

slope of perpendicular line is negative recripocal

OpenStudy (anonymous):

so the new slope is actually -15/3 or -5

OpenStudy (anonymous):

parallel lines have the same slope

OpenStudy (anonymous):

ooo i see

OpenStudy (anonymous):

good :)